OpenStudy (anonymous):

The sum of Tony's age and his older brother's age is 29. The difference between their ages is 5 years. How old is Tony?

terenzreignz (terenzreignz):

First thing you need to do is represent values... do you use systems here?

OpenStudy (anonymous):

Tony = x Tony's elder brother = x + 5 ( You know it's + since he's older ) x + (x+5) = 29 2x = 29-5 2x = 24 x = 12